The packaging of an electric light bulb states that the average length of life of bulbs is 1000 hours. A consumer association thinks that this is an overestimate and tests a random sample of 64 bulbs, recording the life ๐’™ hours, of each bulb. The results are summarized as follows:

โˆ‘ ๐’™ = ๐Ÿ”๐Ÿ‘๐Ÿ—๐Ÿ๐ŸŽ. ๐Ÿ’ and โˆ‘ ๐’™

๐Ÿ = ๐Ÿ”๐Ÿ‘๐Ÿ–๐Ÿ๐Ÿ’๐ŸŽ๐Ÿ”๐Ÿ


1
Expert's answer
2021-03-26T04:32:03-0400

mean "\mu_1" = "\dfrac{\sum x}{n} =\frac{63910.4}{64}=998.6 \ h"

mean "\mu_2" = "\dfrac{\sum x_2}{n} =\frac{63824.061}{64}=997.25 \ h"